OpenStudy (anonymous):

what would the best graph that would represent the solution set.? | 5x -1 | < 1 Need help please..

OpenStudy (anonymous):

Remember this is an absolute value. In piece-wise notation, it's the same as saying: |5x - 1| = 5x - 1 for x >= 0 -(5x - 1) = -5x + 1 for x < 0 So here we have a piece-wise function; which is defined as y = 5x - 1 for x >= 0 and y = -5x + 1 for x < 0. Can you make the graphs of the 2 functions on the appropriate part of the domain?

OpenStudy (anonymous):

I used the definition of an absolute value. It helps someone new to the idea to get the intuitive understanding and get to know exactly what it means. A double-inequality like that would scare anyone who isn't familiar with the idea. When you understand something by definition and intuitively, then the more advanced notation or more difficult concepts related to it come easy.

OpenStudy (anonymous):

yes i needed it as a compound inequality.. its easier to make it into a graph Thanks

OpenStudy (anonymous):

K bro lol.

OpenStudy (anonymous):

Btw, I never noticed the < 1 part -.-. I thought you were just asking how to graph y=|5x - 1|. Oh well...

OpenStudy (anonymous):



OpenStudy (anonymous):

that right.? & lol yeah sorrry @genius12 your response was helpful still

OpenStudy (anonymous):

thanks @Hero